# Loading the stable photon sphere in Weyl conformal gravity metrics

### v1

@@ 1 Introduction
Following the relativity principle proffered by Weyl [15] and Bach [16] at the start of the 20th century, CG was established by Mannheim and Kazanas in 1989 as a promising alternative to GR [17]. While Einstein’s field equations for GR are invariant under coordinate and Lorentz transformations, CG’s field equations possess an additional invariance under conformal (Weyl) transformations, which are local isotropic stretches of the metric preserving angles and shapes [17]:
@@
CG successfully fits flat galaxy rotation curves [18, 19, 20, 21] and addresses the dark energy problem [22] without reliance on the dark sector, as well as being able to reproduce GR results on solar system scales [23]. Furthermore, CG is a promising candidate for QG due to its ghostlessness [24, 25], and it shares conformal invariance with the other forces of the Standard Model [14].
@@
$I_{\rm W}=-\alpha_{g}\int\text{d}^{4}x(-g)^{\frac{1}{2}}C_{\lambda\mu\nu\kappa}C^{\lambda\mu\nu\kappa}.$ (4)
Here, $\alpha_{g}$ is a negative and dimensionless constant associated with the gravitational field [26], while the conformally invariant Weyl tensor $C_{\lambda\mu\nu\kappa}$ is the traceless part of the Riemann tensor $R_{\lambda\mu\nu\kappa}$ [17]:
@@
$W_{\mu\nu}$ is traceless, due to the conformal invariance of this gravitational theory. Subsequently, (6) requires the energy-momentum tensor $T_{\mu\nu}$ to be conformally invariant and thus traceless as well.
@@ 1.1 The Mannheim-Kazanas metric
$\nabla^{4}B(r)=\dfrac{3({T^{0}}_{0}-{T^{r}}_{r})}{4\alpha_{g}B(r)}=\dfrac{-3(\rho+p_{r})}{4\alpha_{g}B(r)}\equiv f(r),$ (12)
@@
where primes denote derivatives in $r$. In the sourceless case when $f(r)=0$, we have a homogeneous fourth-order Poisson equation $\nabla^{4}B(r)=0$, which is solved to give:
@@
$w^{2}=1-6M\gamma-\dfrac{3r^{4}}{4\alpha_{g}}{T^{r}}_{r},$ (17)
where ${T^{r}}_{r}$ is the radial pressure $p_{r}(r)$, which again in the sourceless case is equal to ${T^{0}}_{0}$. The MK parameters in (17) remain constant if $p_{r}\propto 1/r^{4}$, while the vacuum MK parameters of (16) fulfill the third-order constraint of (13) with $4\alpha_{g}{W^{r}}_{r}={T^{r}}_{r}=0$.
@@
Geometric features of nonrotating spacetime metrics include horizons where $B(r)=0$, as well as photon spheres, which are circular null geodesics around the central singularity. These are located at radii where $\mathrm{d}V_{\mathrm{eff}}/\mathrm{d}r=0$, and $V_{\mathrm{eff}}=B(r)/r^{2}$ (see section 2.3), allowing for stable or unstable photon orbits [39]. The GRS metric has one horizon at $r=2\beta$ and one unstable photon sphere at $r=3\beta$. On the other hand, the MK metric has a maximum of three horizons and two photon spheres, with one unstable photon sphere at $r_{\text{ust}}=3\beta$ analogous to the one in GRS, and one stable photon sphere at $r_{\text{st}}=3\beta-2/\gamma$ [40]. In particular, the stability of the latter suggests the likelihood for the buildup of null matter at these radii, which we may refer to as “accumulation points” [40].
In the present work, we solve a toy problem to analyse photon accumulation in CG, by building up an infinitely-thin shell at the radius of the stable photon sphere $r=r_{\mathrm{st}}=3\beta-2/\gamma$. As the stable photon sphere does not exist in the GRS metric, investigating this previously-unexplored phenomenon proves to be an endeavour physically unfeasible in the GRS metric but practicable in CG’s MK metric.
@@ 2.1 Spacetime regions
In the present work, we refer to two types of spacetime regions. In regions where $B(r)>0$, $\mathrm{d}t$ is a spacelike interval, and therefore $t$ acts like a timelike coordinate; hence we refer to these spacetime regions as timelike ($\mathrm{T}$) regions.
@@ 2.2 Horizons
    Event horizon ($\text{H}_{\text{E}}$): For increasing $r$, the spacetime regions are ordered as spacelike to timelike ($\mathrm{S}^{-}\rightarrow\text{H}_{\text{E}}\rightarrow\mathrm{T}$). An S- region interior to $\text{H}_{\text{E}}$ enforces inwards trajectories towards $r=0$.
@@
    Cosmological horizon ($\text{H}_{\Lambda}$): $\mathrm{T}\rightarrow\text{H}_{\Lambda}\rightarrow\mathrm{S}^{+}$ for increasing $r$, like for $\text{H}_{\text{C}}$. $\text{H}_{\Lambda}$, when it exists, is always an outermost horizon [40], which distinguishes itself from $\mathrm{H}_{\mathrm{C}}$. The $\mathrm{S}^{+}$ region exterior to $\text{H}_{\Lambda}$ enforces outwards trajectories to $r\rightarrow+\infty$ [43].
@@ 2.3 Conditions for the existence of photon spheres
A map of these parameters is shown in figure 3, serving as a guideline to selecting appropriate values of $\beta\gamma$ and $\beta^{2}\kappa$. We exclusively select from the yellow domain, corresponding to $\mathrm{T}$ regions. Except for $\beta\gamma\rightarrow$-$\infty$ and $\beta\gamma=2/3$, there is always a negative-enough value of $\beta^{2}\kappa$ that will allow a photon sphere to exist in a T region; crucially, while $\beta^{2}\kappa$ does not affect the radius of the photon sphere, it affects the sign of $B(r_{\mathrm{st}})$ by shifting the potential $V_{\mathrm{eff}}$ up or down, as seen in (23). The region where $r_{\mathrm{st}}\leq 0$ has been blocked out by the dark rectangle; in this region, $r_{\mathrm{st}}>0$ if $\beta<0$. Again, we do not consider such solutions which correspond to negative mass.
@@ 3.1 Thin shell sources
$f_{0}=\dfrac{-3}{4\alpha_{g}}\int\mathrm{d}r\dfrac{\rho(r)+p_{r}(r)}{B(r)}.$ (32)
@@ 3.2 Pressure jump across the shell
$\dfrac{{T^{r}}_{r}}{-\alpha_{g}f_{0}}=\dfrac{2\gamma}{3r_{\mathrm{sh}}}\left(r_{\mathrm{sh}}-r_{\mathrm{st}}\right)\left(r_{\mathrm{sh}}-r_{\mathrm{ust}}\right),$ (41)
@@ 3.3 Photon sphere radii
$=\dfrac{\left(-w_{0}-1\right)\left(\gamma_{0}-\dfrac{(r_{\mathrm{st}}^{-})^{2}f_{0}}{2}\right)+\gamma_{0}\left(w_{0}+\dfrac{(r_{\mathrm{st}}^{-})^{3}f_{0}}{2}+1\right)}{\gamma_{0}\left(\gamma_{0}-\dfrac{(r_{\mathrm{st}}^{-})^{2}f_{0}}{2}\right)}$
$=\dfrac{\dfrac{(r_{\mathrm{st}}^{-})^{2}f_{0}}{2}w_{0}+\dfrac{(r_{\mathrm{st}}^{-})^{2}f_{0}}{2}+\dfrac{(r_{\mathrm{st}}^{-})^{3}f_{0}}{2}\gamma_{0}}{\gamma_{0}\left(\gamma_{0}-\dfrac{(r_{\mathrm{st}}^{-})^{2}f_{0}}{2}\right)}$
@@ 4 Spacetime structures and horizons
Upon loading the photon sphere, $r_{\mathrm{st}}$ does not change; however, the causal structure of the spacetime does indeed change. Plots of $B(r)$ in figure 5 show that $B(r)$ decreases as $f_{0}$ increases, regardless of whether the central singularity is spacelike as $B(r=0)\rightarrow$-$\infty$ (figure 5(a)) or timelike with $B(r=0)\rightarrow+\infty$ (figure 5(b)). Notably, there are no discontinuities in the blackening factor $B(r)$ despite the presence of the shell.
@@ 6 Conclusions and further perspectives
  • Section 3.3 and section 4: Upon loading the stable photon sphere with photons by increasing the amplitude of $f(r)$, its area remains invariant and its radius stays constant, highlighting a previously unseen significance of photon sphere radii in CG’s MK metric [53]. However, locations of horizons may change.

### v2

# Backreactions from loading the stable photon sphere in Weyl conformal gravity

@@ 1 Introduction
Following the relativity principle proffered by Weyl [70] and Bach [7] at the start of the 20th century, CG was established by Mannheim and Kazanas in 1989 as a promising alternative to GR [42]. While Einstein’s field equations for GR are invariant under coordinate and Lorentz transformations, CG’s field equations possess an additional invariance under Weyl transformations, which are local isotropic stretches of the metric preserving angles and shapes [42]:
@@
CG successfully fits flat galaxy rotation curves [21, 24, 53, 54] and addresses the dark energy problem [48] without reliance on the dark sector, as well as being able to reproduce GR results on solar system scales [66]. On the quantum gravity front, CG is a promising QG candidate due to it sharing the conformal invariance of the other forces of the Standard Model [45].
@@
$I_{\rm W}=-\alpha_{\mathrm{g}}\int\text{d}^{4}x(-g)^{\frac{1}{2}}C_{\lambda\mu\nu\kappa}C^{\lambda\mu\nu\kappa}.$ (4)
Here, $\alpha_{\mathrm{g}}$ is a negative and dimensionless constant associated with the gravitational field [46], while the conformally invariant Weyl tensor $C_{\lambda\mu\nu\kappa}$ is the traceless part of the Riemann tensor $R_{\lambda\mu\nu\kappa}$ [42]:
@@
$W_{\mu\nu}$ is traceless, due to the conformal invariance of this gravitational theory. Subsequently, (6) requires the energy-momentum tensor $T_{\mu\nu}$ to be traceless as well.
While higher-order derivative theories usually face scrutiny due to a purported violation of unitarity and the presence of ghosts [3], specifically fourth-order derivative theories, such as CG, have been shown to preserve unitarity and ghostlessness [47, 10]. This proof is attained by assuming that the quantum mechanical Hamiltonian does not need to be Dirac-Hermitian, but is instead $\mathcal{PT}$-symmetric. This allows for the construction of a fourth-order derivative model of the Pais-Uhlenbeck oscillator that is both unitary and ghost-free.
Additionally, a reassessment of the Ostrogradsky instability for higher-order derivative theorems found that the associated ghost is in fact avoidable. This becomes clear when one considers that the original Ostrogradsky theorem only necessarily holds true for Lagrangians of one dynamical variable, while modern field theories have multiple fields [29].
@@ 1.1 The Mannheim-Kazanas metric
$\nabla^{4}B(r)=\dfrac{3({T^{0}}_{0}-{T^{r}}_{r})}{4\alpha_{\mathrm{g}}B(r)}=\dfrac{-3(\rho+p_{r})}{4\alpha_{\mathrm{g}}B(r)}\equiv f(r),$ (12)
@@
where primes denote derivatives in $r$. (12) and (13) are therefore the two independent equations of motion to be solved for the blackening factor in the metric tensor. The metric Ansatz provided in (10) restricts only the diagonal components of the Bach equations to be nonzero, and spherical symmetry imposes that $W^{\phi}_{\phi}=W^{\theta}_{\theta}$; we can always express two of the Bach equations (6) in terms of the other two. For this reason, to generate a blackening factor fulfilling the Bach equations (6), we first solve (12) and impose the constraint from the third-order equation presented in (13).
In the sourceless case when $f(r)=0$, we have a homogeneous fourth-order Poisson equation $\nabla^{4}B(r)=0$, which is solved to give:
@@
$w^{2}=1-6M\gamma-\dfrac{3r^{4}}{4\alpha_{\mathrm{g}}}{T^{r}}_{r},$ (17)
where ${T^{r}}_{r}$ is the radial pressure $p_{r}(r)$, which again in the sourceless case is equal to ${T^{0}}_{0}$. The MK parameters in (17) remain constant if $p_{r}\propto 1/r^{4}$, while the vacuum MK parameters of (16) fulfill the third-order constraint of (13) with $4\alpha_{\mathrm{g}}{W^{r}}_{r}={T^{r}}_{r}=0$.
@@
Geometric features of nonrotating spacetime metrics include horizons where $B(r)=0$, as well as photon spheres, which are circular null geodesics around the central singularity. These are located at radii where $\mathrm{d}V_{\mathrm{eff}}/\mathrm{d}r=0$, and $V_{\mathrm{eff}}=B(r)/r^{2}$ (see section 2.3), allowing for stable or unstable photon orbits [57]. The GRS metric has one horizon at $r=2\beta$ and one unstable photon sphere at $r=3\beta$. On the other hand, the MK metric has a maximum of three horizons and two photon spheres, with one unstable photon sphere at $r_{\text{ust}}=3\beta$ analogous to the one in GRS, and one stable photon sphere at $r_{\text{st}}=3\beta-2/\gamma$ [67]. In particular, the stability of the latter suggests the likelihood for the buildup of null matter at these radii, which we may refer to as “accumulation points” [67]. As proven in [17], photon spheres always come in pairs; naturally, as $\gamma=0$ in GRS, the stable photon sphere in this metric can be thought to be located at $r\rightarrow+\infty$.
In the present work, we solve a toy problem to analyse photon accumulation in CG, by building up an infinitely-thin shell at the radius of the stable photon sphere $r=r_{\mathrm{st}}=3\beta-2/\gamma$. Previous works have attempted the investigation of self-gravitating photons at the photon sphere in the GRS metric [22]; investigating this previously-unexplored phenomenon with the MK metric proves to be more physically feasible, as the photons are not required to teeter exactly on the unstable extremum of the effective potential. As has also been seen in similar analyses of GR metrics [22], such an investigation would prove to elucidate the stability of the MK metric when subject to backreactions from accumulating null matter.
@@ 2.1 Spacetime regions
In the present work, we refer to two types of spacetime regions. In regions where $B(r)>0$, $\mathrm{d}t$ is a timelike interval, and therefore $t$ acts like a timelike coordinate; hence we refer to these spacetime regions as timelike ($\mathrm{T}$) regions.
@@ 2.2 Horizons
    Event horizon ($\text{H}_{\text{E}}$): For increasing $r$, the spacetime regions are ordered as spacelike to timelike ($\mathrm{S}^{-}\rightarrow\text{H}_{\text{E}}\rightarrow\mathrm{T}$). An S- region interior to $\text{H}_{\text{E}}$ enforces inward trajectories towards $r=0$.
@@
    Cosmological horizon ($\text{H}_{\Lambda}$): $\mathrm{T}\rightarrow\text{H}_{\Lambda}\rightarrow\mathrm{S}^{+}$ for increasing $r$, like for $\text{H}_{\text{C}}$. $\text{H}_{\Lambda}$, when it exists, is always an outermost horizon [67], which distinguishes itself from $\mathrm{H}_{\mathrm{C}}$. The $\mathrm{S}^{+}$ region exterior to $\text{H}_{\Lambda}$ enforces outward trajectories to $r\rightarrow+\infty$ [51].
@@ 2.3 Conditions for the existence of photon spheres
A map of these parameters is shown in figure 3, serving as a guideline to selecting appropriate values of $\beta\gamma$ and $\beta^{2}\kappa$, such that physically meaningful stable photon sphere radii may be obtained. We exclusively select from the yellow domain, corresponding to $\mathrm{T}$ regions. Except for $\beta\gamma\rightarrow$-$\infty$ and $\beta\gamma=2/3$, there is always a negative-enough value of $\beta^{2}\kappa$ that will allow a photon sphere to exist in a T region; crucially, while $\beta^{2}\kappa$ does not affect the radius of the photon sphere, it affects the sign of $B(r_{\mathrm{st}})$ by shifting the potential $V_{\mathrm{eff}}$ up or down, as seen in (23). The region where $r_{\mathrm{st}}\leq 0$ has been blocked out by the dark rectangle; in this region, $r_{\mathrm{st}}>0$ if $\beta<0$. Again, we do not consider such solutions which correspond to negative mass.
@@
An analysis of the topology of stationary black holes in asymptotically flat spacetimes in both GR and modified gravity theories has shown the guaranteed presence of at least one photon sphere radius in such spacetimes [18]. While the metric we are considering here is not asymptotically flat due to the long-range linear and quadratic terms, the aforementioned work does not preclude the existence of photon spheres in non-asymptotically flat spacetimes. For instance, stationary GR spacetimes with de Sitter and Anti-de Sitter backgrounds still possess photon spheres [15, 72].
@@ 3.1 Thin shell sources
$f_{0}=\dfrac{-3}{4\alpha_{\mathrm{g}}}\int\mathrm{d}r\dfrac{\rho(r)+p_{r}(r)}{B(r)}.$ (32)
@@ 3.2 Pressure jump across the shell
$\dfrac{{T^{r}}_{r}}{-\alpha_{\mathrm{g}}f_{0}}=\dfrac{2\gamma}{3r_{\mathrm{sh}}}\left(r_{\mathrm{sh}}-r_{\mathrm{st}}\right)\left(r_{\mathrm{sh}}-r_{\mathrm{ust}}\right),$ (41)
@@ 3.3 Photon sphere radii
We note several discrepancies between our CG results, versus those found in similar studies conducted in GR.
First, we note that unlike what is found in [22], the accumulation of photons at the stable photon sphere of the MK metric does not give rise to additional spherical orbits. In our case, this may be justified by a simple algebraic argument: the interior ($r<r_{\mathrm{st}}$) and exterior ($r>r_{\mathrm{st}}$) effective potentials $V_{\mathrm{eff}}$ generated by the MK metric together form a smooth cubic function in $\beta/r$, which at most have one unstable maximum and one stable minimum. The unstable maximum will not move when $r_{\mathrm{ust}}<r_{\mathrm{st}}$ and will move if $r_{\mathrm{ust}}>r_{\mathrm{st}}$, but this in principle has no effect on the total number of photon spheres obtainable in the metric.
Second, [19] finds that accumulating photons at the stable photon sphere of an ultracompact continuum-shell star in GR serve to push the photon sphere towards smaller $r$. Again, this may be justified by a purely algebraic argument: the presence of a linear $\gamma\,r$ in the blackening factor gives rise to the stable photon sphere radius $r_{\mathrm{st}}=(-1-w)/\gamma$. $w$ jumps up across the shell but $\gamma$ drops down, and the total change in $r_{\mathrm{st}}$ amounts to zero; in the absence of $\gamma$, such as in e.g. GRS, the sole contribution is from $w$, and the overall effect is to push the photon sphere inwards. Therefore, the contrast in the effect of photon accumulation between [19] and the present work may be an indication that CG spacetimes are generally less sensitive to perturbations and less prone to instability than GR spacetimes.
@@ 4 Spacetime structures and horizons
Upon loading the photon sphere, $r_{\mathrm{st}}$ does not change; however, the causal structure of the spacetime may indeed change. Plots of $B(r)$ in figure 5 show that $B(r)$ decreases as $f_{0}$ increases, regardless of whether the central singularity is spacelike as $B(r=0)\rightarrow$-$\infty$ (figure 5(a)) or timelike with $B(r=0)\rightarrow+\infty$ (figure 5(b)). Notably, there are no discontinuities in the blackening factor $B(r)$ despite the presence of the shell.
@@ 6 Conclusions and further perspectives
  • Section 3.3 and section 4: Upon loading the stable photon sphere with photons by increasing the amplitude of $f(r)$, its area remains invariant and its radius stays constant. This highlights a previously unseen significance of photon sphere radii in CG’s MK metric [49] and the increased stability of the Mannheim-Kazanas metric in comparison to some metrics in GR [19]. However, locations of horizons may change.